OpenStudy (anonymous):

can someone help me out with this problem: x^(2/3) + x^(1/3) - 6 =0

OpenStudy (anonymous):

Im asked to find x

OpenStudy (anonymous):

already found the answer

OpenStudy (anonymous):

Consider x^1/3 = y ---- [1] it will form a quadratic equation as y^2 - y + 6 = 0 solve y and the put the value of y in [1]

OpenStudy (anonymous):

sorry it will form a quadratic equation as y^2 + y 6 6 = 0

OpenStudy (anonymous):

y^2 + y - 6 = 0

OpenStudy (anonymous):

Im rising every element in the equation to the third power, so x2 + x -216=0 Am I right?